# Equilibrium Points and Surface Dynamics About Comet 67P/Churyumov-Gerasimenko

### v1

@@ 3.7 Equilibrium Points and Zero-Velocity Curves
We identified five equilibrium points based on an adequate 3-D polyhedral shape model of comet 67P, which comprises 96,834 faces and 48,420 vertices [17]. Figure 7 shows the location of the five equilibrium points with different topological characteristics viewed in the $xOy$ plane. The equilibrium points E1 and E3 are unstable and classified as Case II [33]. Equilibrium point E4 is also unstable and classified as Case V. Unlike Scheeres [16], which identified four equilibrium points around comet 67P, all of them unstable, the equilibrium points E2 and E5 were linearly stable and classified as Case I, indicating that particles surrounding these equilibrium points may have timescales greater than the unstable equilibrium points E1, E3, and E4. These equilibrium point regions may offer reduced fuel costs under idealized conditions. Additionally, equilibrium point E5 is located inside the comet 67P close to its center of mass.
@@
The equilibrium points pair E1-E5 has a large $z$-component ($\sim 10^{-1}$) when compared with the others equilibrium points ($\sim 10^{-2}$). Table 2 also illustrates that the radial distance to the center of mass highlights an axial symmetry that occurs particularly between the equilibrium point pairs E1–E3 and E2–E4, which are located on opposite sides of the comet 67P. Furthermore, the equilibrium points E1 and E3 lie above the $xOy$ plane, while the equilibrium points E2, E4, and E5 remain below it.
@@ 4.2 Solar Radiation Pressure Perturbation
where $\mathbf{r}_{S}$ is the position vector of the Sun relative to 67P, $\|\mathbf{r}_{S}-\mathbf{r}\|$ is the distance from the particle’s located at the barycenter of each face to the Sun, $S_{0}$ is a solar constant or radiation flux density at the distance of the astronomical unit $R_{0}$} (AU), $c$ is the speed of light, and $Q_{pr}$ is the dimensionless efficiency factor for radiation pressure, which depends on the properties (e.g., density, shape, size) of the particle. In the mathematical model, this value is assumed to be $1$ to represent the value of an ideal material. The values of the adopted SRP parameters are presented in Tab. 3. Furthermore, the particles in the mathematical model were assumed to be fully spherical, in order to vary the area-to-mass ratio, as shown in Eq. (9):
@@ 4.3 Comparison Between Dipole Segment and Polyhedron Models
where $U_{DS}(x,y)$ is the gravitational potential from the DS model and $U_{p}(x,y)$ is the gravitational potential from the polyhedron model, both computed in points around comet 67P, in the $xOy$ plane.
@@
Figure 10: Gravitational potential relative error between Dipole Segment (DS) and polyhedron models. The colorbox indicates the relative error (%) around comet 67P for points in the $xOy$ plane.
@@ 5 Evolution of the Characteristic Curves in the Dipole Segment Model Approach
Table 5 presents the topological classification of the characteristic curves that enables the identification of characteristics of orbital families without the need for direct visualization of the orbits [38]. It shows the common features shared by all orbits within a family, as well as the variations occurring at the endpoints of each curve, based on the points encircled by the orbit, namely, the central body (denoted by $DS$) and the equilibrium points, whose locations and designations correspond to those in Fig. 7, with $L_{2}$ and $L_{4}$ being the triangular points collectively represented by the single letter $T$.
@@ 5.1 Stability of the Symmetric Periodic Orbits
Regarding vertical stability, which defines the stability outside the plane of the planar orbit, note that in the Families $1$, 55 and $12$ there are stable and unstable orbits; in Families $2$, 44 and $11$ only stability; and in Family $3$ the maximum index is $S_{v}=2.000398$ (Fig. 16).
@@
If $0<S_{v}<2$ and $S_{v}/2=|cos(2\pi m/n)$ with $m,n\in\mathbb{Z}$ and $m<n$, then the SPO family can bifurcate into a 3D family of periodic orbits with multiplicity $nM$.
@@
- The determinant of the monodromy matrix must be equal to $1$;
- If the eigenvalues are complex conjugates of magnitude $1$, the system is stable; if they are real reciprocals, the system is unstable.
- For systems with a monodromy matrix $\mathbb{R}^{2\times 2}$ and a determinant equal to $1$, $S_{v}=\mid\lambda_{v1}+\lambda_{v2}\mid=\mid\lambda_{v1}+\frac{1}{\lambda_{v1}}\mid=\mid\operatorname{tr}{(\Phi_{v}(T))\mid}$, since $\operatorname{tr}$ is the trace of the monodromy matrix .

### v2

# arXiv:2509.11613v2[astro-ph.EP] 01 Mar 2026

@@ 3.7 Equilibrium Points and Zero-Velocity Curves
We identified five equilibrium points based on an adequate 3-D polyhedral shape model of comet 67P, which comprises 96,834 faces and 48,420 vertices [17]. Figure 7 shows the location of the five equilibrium points with different topological characteristics viewed in the xOyxOy plane. The equilibrium points E1 and E3 are unstable and classified as Case II [33]. Equilibrium point E4 is also unstable and classified as Case V. Unlike Scheeres [16], which identified four equilibrium points around comet 67P, all of them unstable, the equilibrium points E2 and E5 were linearly stable and classified as Case I, indicating that particles surrounding these equilibrium points may have timescales greater than the unstable equilibrium points E1, E3, and E4. These equilibrium point regions may offer reduced fuel costs under idealized conditions. Additionally, equilibrium point E5 is located inside the comet 67P close to its center of mass.
@@
The equilibrium points pair E1-E5 has a large $z$-component ($\sim 10^{-1}$) when compared with the others equilibrium points ($\sim 10^{-2}$). Table 2 also illustrates that the radial distance to the center of mass highlights an axial symmetry that occurs particularly between the equilibrium point pairs E1–E3 and E2–E4, which are located on opposite sides of the comet 67P. Furthermore, the equilibrium points E1 and E3 lie above the xOyxOy plane, while the equilibrium points E2, E4, and E5 remain below it.
@@ 4.2 Solar Radiation Pressure Perturbation
where $\mathbf{r}_{S}$ is the position vector of the Sun relative to 67P, $\|\mathbf{r}_{S}-\mathbf{r}\|$ is the distance from the particle’s located at the barycenter of each face to the Sun, $S_{0}$ is a solar constant or radiation flux density at the distance of the astronomical unit $R_{0}$} (AU), $c$ is the speed of light, and $Q_{pr}$ is the dimensionless efficiency factor for radiation pressure, which depends on the properties (e.g., density, shape, size) of the particle. In the mathematical model, this value is assumed to be 11 to represent the value of an ideal material. The values of the adopted SRP parameters are presented in Tab. 3. Furthermore, the particles in the mathematical model were assumed to be fully spherical, in order to vary the area-to-mass ratio, as shown in Eq. (9):
@@ 4.3 Comparison Between Dipole Segment and Polyhedron Models
where $U_{DS}(x,y)$ is the gravitational potential from the DS model and $U_{p}(x,y)$ is the gravitational potential from the polyhedron model, both computed in points around comet 67P, in the xOyxOy plane.
@@
Figure 10: Gravitational potential relative error between Dipole Segment (DS) and polyhedron models. The colorbox indicates the relative error (%) around comet 67P for points in the xOyxOy plane.
@@ 5 Evolution of the Characteristic Curves in the Dipole Segment Model Approach
Table 5 presents the topological classification of the characteristic curves that enables the identification of characteristics of orbital families without the need for direct visualization of the orbits [38]. It shows the common features shared by all orbits within a family, as well as the variations occurring at the endpoints of each curve, based on the points encircled by the orbit, namely, the central body (denoted by DSDS) and the equilibrium points, whose locations and designations correspond to those in Fig. 7, with $L_{2}$ and $L_{4}$ being the triangular points collectively represented by the single letter $T$.
@@ 5.1 Stability of the Symmetric Periodic Orbits
Regarding vertical stability, which defines the stability outside the plane of the planar orbit, note that in the Families 11, 55 and $12$ there are stable and unstable orbits; in Families 22, 44 and $11$ only stability; and in Family 33 the maximum index is $S_{v}=2.000398$ (Fig. 16).
@@
If $0<S_{v}<2$ and $S_{v}/2=|cos(2\pi m/n)$ with $m,n\in\mathbb{Z}$ and $m<n$, then the SPO family can bifurcate into a 3D family of periodic orbits with multiplicity nMnM.
@@
- The determinant of the monodromy matrix must be equal to 11;
- If the eigenvalues are complex conjugates of magnitude 11, the system is stable; if they are real reciprocals, the system is unstable.
- For systems with a monodromy matrix $\mathbb{R}^{2\times 2}$ and a determinant equal to 11, $S_{v}=\mid\lambda_{v1}+\lambda_{v2}\mid=\mid\lambda_{v1}+\frac{1}{\lambda_{v1}}\mid=\mid\operatorname{tr}{(\Phi_{v}(T))\mid}$, since $\operatorname{tr}$ is the trace of the monodromy matrix .